OpenStudy (anonymous):

Find a function g(x) such that g(g(x)) = 6x − 8. Need help..

OpenStudy (anonymous):

Let's suppose \(g(x)\) is a line. Then it has for form: \[ g(x) = Ax+B \]

OpenStudy (anonymous):

Note: I'm not sure that \(g(x)\) is a line, it's just an educated guess for now.

OpenStudy (anonymous):

In fact, I'm not even sure \(g(x)\) has a single solution. It may have many.

OpenStudy (anonymous):

Anyway, the assumption would give us: \[ g(g(x)) = A(Ax+B)+B = A^2x + AB + B \]

OpenStudy (anonymous):

So we have: \[ A^2x + AB + B = 6x-8 \\ (A^2)x + (AB + B) = (6)x+(-8) \\ A^2 = 6, \quad AB + B = -8 \]Think you can solve for \(A\) and \(B\)? @rashley284

OpenStudy (anonymous):

but what is g(x)?

OpenStudy (anonymous):

We said\[ g(x) = Ax+B \]So if you solve for \(A\) and \(B\), then you have found \(g(x)\).

OpenStudy (anonymous):

g(g(x))=6x-8 . I don't get it.

OpenStudy (anonymous):

@rashley284 Okay, I'll try to explain better. We're going to assume that \(g(x)\) could be a line. If it is a line, then it has the following form: \[ y = mx+b \]Does this make sense so far?

OpenStudy (anonymous):

If we assume that \(g(x)\) is a line, then we need to find it's slope and intercept. \[ g(x) = mx+b \]For now, we will just keep them as variables \(m\) and \(b\).

OpenStudy (anonymous):

@rashley284 Do you follow?

OpenStudy (anonymous):

yes. . i get it now. Thank you.

OpenStudy (anonymous):

Do I need to explain it anymore? What answer are you getting?

OpenStudy (anonymous):

so the slope is 6 and the y-intercept is (0,-8)? @wio